A secretary makes 2 errors per page on the average. What is the probability that on the next page she makes


a) 4 or more errors?


b) No errors?


1
Expert's answer
2022-04-21T13:16:26-0400

According to the Poisson's distribution "P(x=k)=\frac{\lambda^ke^{-\lambda}}{k!},"

where "\lambda" - average value


"a)~P(x\ge4)=1-P(x\le3)=\\
1-(P(x=0)+P(x=1)+\\
+P(x=2)+P(x=3))=\\
=1-(\frac{2^0}{0!}+\frac{2^1}{1!}+\frac{2^2}{2!}+\frac{2^3}{3!})/e^2=\\
=1-(1+2+2+\frac{4}{3})/e^2\approx0.143=14.3\%"


"b)~P(x=0)=\frac{2^0}{e^20!}=1/e^2\approx0.135=13.5\%"